I need help with this problem
x^(3/2) - 2x^(3/4) +1 =0

Answer by xinxin(76)

Assume x^(3/4)=A, then x^(3/2)=x^(3/4*2)=A^2. The function turns out to be A^2-2A+1=0
Get A=1, therefore x=1